# Resumming Post-Minkowskian and Post-Newtonian gravitational waveform expansions

### v1

@@ 1 Introduction
Recent advances in the sensitivity of experimental apparatus for detecting gravitational waves (GWs) necessitate increasingly precise theoretical computations. The next generation of space-based experiments, such as the Laser Interferometer Space Antenna (LISA), promises to revolutionize the GW detection by significantly enhancing sensitivity and exploring entirely new wavelength ranges compared to ground-based interferometers LISA:2022kgy . LISA will detect long-wavelength GWs originating from sources like supermassive black hole mergers, extreme mass-ratio inspirals (EMRIs), galactic binaries, and the Stochastic Gravitational Wave Background—sources largely inaccessible to ground-based observatories.
@@
The gravitational signals from these sources are diverse. For instance, GWs from supermassive black hole mergers will produce strong signals with large signal-to-noise ratios, while those from EMRIs, though weaker, will carry exceptionally detailed information about the gravitational sources. Indeed, EMRIs waveforms, spanning thousands of cycles during their inspiral phase, provide a unique opportunity to probe fundamental questions about the nature of gravitational sources, including their compactness, the presence or absence of horizons, light rings, echoes Cardoso:2019rvt , and the validity of Einstein’s theory of gravity and its possible extensions Gair:2012nm ; Yunes:2013dva .
@@
Two main analytical tools have been developed for studying the problem we just described: the multipole expansion Thorne:1980ru ; Damour:1982wm ; Blanchet:2013haa and perturbation theory Teukolsky:1973ha ; Poisson:1993vp ; Tagoshi:1993dm ; Poisson:1994yf ; Shibata:1994jx ; Tagoshi:1994sm ; Tanaka:1996lfd ; Mano:1996vt ; Mino:1997bx ; Suzuki:1998vy ; Fujita:2010xj ; Fujita:2011zk . The latter, which applies to EMRI systems, is the focus of this paper. In this framework, GWs are computed by solving the Einstein equations to linear order in the perturbations induced by the motion of a lighter body orbiting around a very massive black hole (BH). The radial and angular propagations of the waves are governed by ordinary differential equations of the confluent Heun type. Analytic solutions of these equations are not known, but it is possible to write them as an expansion in the Newton constant (Post-Minkowskian, or PM, approximation) or in the orbital velocity (Post-Newtonian, or PN, approximation).
@@
In the case of circular orbits, the PM and PN approximations are linked and characterised by a single expansion parameter $v$ (the tangential velocity). The current state-of-the-art for the emitted energy in perturbation theory is the 22nd PN (order $v^{44}$) computation in Fujita:2012cm . The expansion is known to be converging slowly Cutler:1993vq ; Poisson:1993vp due to the presence of logarithmic terms in the velocity, see also Damour:2008gu ; Damour:2009kr ; Buonanno:2009zt ; Damour:2012ky . In Aminov:2020yma ; Bianchi:2021xpr ; Bonelli:2021uvf ; Bonelli:2022ten ; Consoli:2022eey ; Bautista:2023sdf ; Aminov:2023jve ; ToVSapQNM ; Fucito:2023afe novel techniques, capitalizing on recent advances in quantum field theories, have been exploited to derive an exact formula for the Heun connection matrix and alternative formulations of the waveforms were proposed in terms of the instanton partition function of a $SU(2)^{2}$ quiver gauge theory. Following this path the infinite tower of logarithmic terms (tails and tails of tails) have been resummed into exponentials.
@@
More recently, multipole Bini:2023fiz ; Brandhuber:2023hhy ; DeAngelis:2023lvf ; Brandhuber:2023hhl ; Herderschee:2023fxh ; Georgoudis:2023eke ; Georgoudis:2023lgf ; Georgoudis:2023ozp ; Elkhidir:2023dco ; Bini:2024ijq ; Bini:2024rsy ; Georgoudis:2024pdz ; Alessio:2024onn ; Brunello:2024ibk and perturbation theory Fucito:2024wlg methods have been applied to the study of scattering processes (open trajectories) in gravity Kovacs:1977uw ; Kovacs:1978eu , a field where we can largely benefit from the impressive developments and refinements of scattering amplitude techniques in the last few years (see DiVecchia:2023frv for a review). These systems provide a richer set up, where the couplings of the PM and PN expansions are no longer linked to each other. For example, in the scattering amplitude language, the PM expansion corresponds to the expansion in loops. At a given order in the coupling, the amplitude is exact in the velocity, i.e. it resums the infinite PN series. To reproduce this result in perturbation theory, one has to resum the double expansion in one of the two variables.
@@
For simplicity we focus on the geodesic motion in a Schwarzschild geometry and parametrize the mass, wavefrequency and radial distance by the two dimensionless combinations $z=2M/r$ and $y=2i\omega r$, assuming $x=4{\rm i}\omega M$ is small. We will show that, at order $x^{k}$, the dependence of the waveform on the variable $y$ or $z$ can be resummed and written in terms of a single hypergeometric function and its derivative, with coefficients given by two polynomials of degree $k$, that can be determined recursively. The two limits $x\to 0$ and $y$ or $z$ finite lead to two alternative hypergeometric representations of the confluent Heun function, describing the PM and PN approximations respectively. The parameters of the hypergeometric function and the coefficients of the polynomials are all expressed in terms of a characteristic function $a(\ell)$ that, in the MST Tagoshi:1993dm ; Poisson:1994yf ; Shibata:1994jx ; Tagoshi:1994sm ; Tanaka:1996lfd ; Mano:1996vt ; Mino:1997bx ; Suzuki:1998vy ; Fujita:2010xj ; Fujita:2011zk formulation of the confluent Heun solution, is the "renormalized angular momentum", while in the description of this paper Aminov:2020yma ; Bianchi:2021xpr ; Bonelli:2021uvf ; Bonelli:2022ten ; Consoli:2022eey ; Bautista:2023sdf ; Aminov:2023jve ; ToVSapQNM ; Fucito:2023afe ; Fucito:2024wlg , it is the Seiberg-Witten (SW) period of the ${\cal N}=2$ supersymmetric gauge theory. In this framework the gauge couplings of the quiver gauge theory are connected to the PM and PN expansion parameters as $q_{1}=z$ and $q_{2}=y$. The same techniques apply to D-branes, topological stars and other gravitational backgrounds governed by an equation of the Heun type. See Bianchi:2024rod ; Bianchi:2024vmi for a recent application to the case of topological stars.
@@
To keep the material self-contained, in this paper we present a derivation of the Heun solutions from first principles with no reference to supersymmetric gauge theory or localization, which sit in the background as inspiration to some of our ansätze, that nonetheless could be also accepted as such. These results provide a first principle proof of the Heun connection formula derived in Aminov:2020yma ; Bianchi:2021xpr ; Bonelli:2021uvf ; Bonelli:2022ten from the crossing symmetry of the conformal blocks of a conformal field theory (CFT) . Our results provide an algorithm which is very efficient when implemented in a Mathematica code and that will allow us to compute the gravitational waveform in the case of circular orbits up to the 30th PN order. This latter result will be tested against a numerical solution and the 22nd PN order computation available in literature Fujita:2012cm . Numerical studies of unbounded systems are postponed to future investigations.
@@
This is the plan of the paper: in Section 2 we derive the PM and PN expansion of the confluent Heun function. In Section 3 we apply the results of Section 2 to the study of the gravitational waveforms emitted by a particle moving in a Schwarzschild background. In Section 4 we test our results against those obtained via the MST method and numerical integration. In Section 5 we extend the results to include the case of the non-confluent Heun function. Section 6 contains our conclusions.
2 The confluent Heun equation
@@
$\beta$ $=$ $-x(\kappa_{1}+\kappa_{3})\qquad,\qquad\delta=\kappa_{2}+\kappa_{3}\qquad,\qquad\gamma=1-\kappa_{2}+\kappa_{3}$
$\zeta$ $=$ $-x\qquad,\qquad\qquad\qquad q=u-\kappa_{3}^{2}-x(\kappa_{1}+\kappa_{3})$ (3)
@@
$G_{\alpha}(z)\underset{x,z\to 0\atop x/z~{}{\rm finite}}{\approx}H^{0}_{\alpha}({\textstyle\frac{x}{z}})\qquad,\qquad G_{\alpha}(z)\underset{x\to 0\atop z~{}{\rm finite}}{\approx}H^{1}_{\alpha}(z)$ (7)
@@
$H^{0}_{\alpha}(y)$ $=$ $y^{-\alpha a+\kappa_{1}}{}_{1}F_{1}(\kappa_{1}{-}\alpha a,1{-}2\alpha a,y)$
@@
${\rm Near~{}zone}:\qquad\qquad G_{\alpha}(z)=z^{\alpha\,a{-}\kappa_{1}}\sum_{i,j=0}^{\infty}G_{\alpha ij}(a)\,z^{i}\,({\textstyle\frac{x}{z}})^{j}\qquad x\ll z\ll 1$ (11)
with $a(u)$ given in (9). The quantity $a$ plays the role of a ‘Floquet exponent’, specifying the monodromy of $G_{\alpha}$ around a contour enclosing the points 0 and $x$. As we will see, the solutions of (4) and the connection formulae take a simpler form if written in terms of the variable $a$ rather than $u$. So we write
@@
2.1 PM approximation: exterior region
@@
$P_{0}(y)=1+\sum_{i=1}^{k}\sum_{j=0}^{i-1}c_{ij}\,x^{i}y^{j-i}\qquad,\qquad\widehat{P}_{0}(y)=\sum_{i=1}^{k}\sum_{j=0}^{i-1}\widehat{c}_{ij}\,x^{i}y^{j-i}\$ (15)
are polynomials of order $k$ in $x$ [^2] For $k=1$ M.Billò, M.L.Frau and A.Lerda first obtained (14) resumming the contributions of the Young tableaux entering the partition function of the supersymmetric gauge theory.. We plug this ansatz into (4) and use the differential equation to express higher derivatives of $H^{0}_{\alpha}$ in terms of $H^{0}_{\alpha}$ and its first derivative themselves. Therefore one finds a system of algebraic linear equations that determines the coefficients $c_{ij}$, $\widehat{c}_{ij}$ and $u_{i}$ (whose complete forms are given in (71)). Setting $c_{00}=1$, one finds, for the first term in the recursion
@@
$c_{i,i-1}=u_{i}\qquad,\qquad\gamma_{\kappa}=1+\sum_{i=1}^{\infty}\widehat{c}_{i,i-1}x^{i}={\rm exp}\left\{-\sum_{i=1}^{\infty}{\partial_{\kappa_{1}}u_{i}\over i}x^{i}\right\}$ (17)
@@
The subscript "inst" reminds the reader that this quantity is interpreted as the instanton prepotential in the gauge theory description of the Heun equation. Indeed in the limit $z\to 0$, where one of the two gauge couplings is turned off, the quiver theory reduces to a ${\cal N}=2$ supersymmetric $SU(2)$ gauge theory with three massive fundamental flavours with prepotential ${\cal F}_{\rm inst}$.
@@
2.2 PN approximation: interior region
@@
Higher orders follow from the ansatz
@@
$P_{1}(z)=1+\sum_{j=1}^{k}\sum_{i=0}^{j-1}d_{ij}\,z^{i-j}x^{j}\qquad,\qquad\widehat{P}_{1}(z)=\sum_{j=1}^{k}\sum_{i=0}^{j}\widehat{d}_{ij}\,z^{i-j}x^{j}\$
@@
$d_{01}=\frac{2\left(a^{2}-\kappa_{1}^{2}\right)}{4a^{2}-1}\qquad,\qquad\widehat{d}_{11}=-\widehat{d}_{01}=\frac{\left(2\kappa_{1}-1\right)}{4a^{2}-1}$ (26)
@@
$\widetilde{H}^{1}_{\alpha}(z)=z^{a-\kappa_{1}}(1{-}z)^{{1{+}\alpha\over 2}(1{-}\kappa_{2}{-}\kappa_{3})}\,_{2}F_{1}\left({\textstyle\frac{1{+}\alpha}{2}}{-}\alpha\kappa_{2}{+}a,{\textstyle\frac{1{+}\alpha}{2}}{-}\alpha\kappa_{3}{+}a;1{+}\alpha\left(1{-}\kappa_{2}{-}\kappa_{3}\right);1{-}z\right)$
@@
2.3 Boundary conditions and connection formulae
@@
3 Black hole perturbation
@@
3.1 Heun gravity correspondence
@@
with $\Delta=r(r-2M)$. We are interested in the $\psi_{4}$ fundamental mode Newman:1961qr ; Teukolsky:1973ha corresponding to $s=-2$. (38) can be put into the CHE form (4) after the identifications (look at Appendix A for details)
@@
$R_{\rm in}(z)$$\approx$ $\left\{\begin{array}[]{lll}B^{\rm in}_{-}e^{-{x\over 2z}}z^{\kappa_{1}-{3\over 2}}+B^{\rm in}_{+}e^{{x\over 2z}}z^{-\kappa_{1}-{1\over 2}}&{}{}{}{}{}{}{}&\quad z\to 0\\ D^{\rm in}_{-}(1-z)^{1+\kappa_{2}+\kappa_{3}\over 2}&&\quad z\to 1\\ \end{array}\right.$ (42)
$R_{\rm up}(z)$$\approx$ $\left\{\begin{array}[]{lll}B^{\rm up}_{+}e^{{x\over 2z}}z^{-\kappa_{1}-{1\over 2}}&{}{}{}{}{}{}{}&\quad z\to 0\\ D^{\rm up}_{-}(1-z)^{1+\kappa_{2}+\kappa_{3}\over 2}+D^{\rm up}_{+}(1-z)^{3-\kappa_{2}-\kappa_{3}\over 2}&&\quad z\to 1\\ \end{array}\right.$ (45)
@@
$G(r,r^{\prime})={1\over W}\left\{\begin{array}[]{ccc}R_{\rm in}(r^{\prime})R_{\rm up}(r)&{}{}{}{}{}{}{}\hfil&r^{\prime}<r\\ R_{\rm in}(r)R_{\rm up}(r^{\prime})&{}{}{}{}{}{}{}\hfil&r<r^{\prime}\\ \end{array}\right.$ (46)
@@
$R_{3}(a,y)=\frac{(7-2a)y^{3}}{8(2a-1)(2a-3)(a-1)}+\frac{(11-6a)x}{4(2a-1)}$ (54)
$R_{4}(a,y)=\frac{\left(4a^{2}{-}72a{+}163\right)y^{4}}{512(a{-}2)(a{-}1)(2a{-}3)(2a{-}1)}+\frac{\left({-}4a^{2}{+}48a{-}119\right)yx}{64(2a{-}1)(a{-}1)}{+}\frac{\left(8a^{3}{-}4a^{2}{-}18a{+}9\right)x^{2}}{64y^{2}(a{+}1)}$

### v2

@@ I Introduction
Recent advances in the sensitivity of experimental apparatus for detecting gravitational waves (GWs) necessitate increasingly precise theoretical computations. The next generation of space-based experiments, such as the Laser Interferometer Space Antenna (LISA), promises to revolutionize the GW detection by significantly enhancing sensitivity and exploring entirely new wavelength ranges compared to ground-based interferometers [1]. LISA will detect long-wavelength GWs originating from sources like supermassive black hole mergers, extreme mass-ratio inspirals (EMRIs), galactic binaries, and the Stochastic Gravitational Wave Background—sources largely inaccessible to ground-based observatories.
@@
The gravitational signals from these sources are diverse. For instance, GWs from supermassive black hole mergers will produce strong signals with large signal-to-noise ratios, while those from EMRIs, though weaker, will carry exceptionally detailed information about the gravitational sources. Indeed, EMRIs waveforms, spanning thousands of cycles during their inspiral phase, provide a unique opportunity to probe fundamental questions about the nature of gravitational sources, including their compactness, the presence or absence of horizons, light rings, echoes [2], and the validity of Einstein’s theory of gravity and its possible extensions [3, 4].
@@
Two main analytical tools have been developed for studying the problem we just described: the multipole expansion [5, 6, 7] and perturbation theory [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. The latter, which applies to EMRI systems, is the focus of this paper. In this framework, GWs are computed by solving the Einstein equations to linear order in the perturbations induced by the motion of a lighter body orbiting around a very massive black hole (BH). The radial and angular propagations of the waves are governed by ordinary differential equations of the confluent Heun type. Analytic solutions of these equations are not known, but it is possible to write them as an expansion in the Newton constant (Post-Minkowskian, or PM, approximation) or in the orbital velocity (Post-Newtonian, or PN, approximation).
@@
In the case of circular orbits, the PM and PN approximations are linked and characterised by a single expansion parameter $v$ (the tangential velocity). The current state-of-the-art for the emitted energy in perturbation theory is the 22nd PN (order $v^{44}$) computation in [20]. The expansion is known to be converging slowly [21, 9] due to the presence of logarithmic terms in the velocity, see also [22, 23, 24, 25]. In [26, 27, 28, 29, 30, 31, 32, 33, 34] novel techniques, capitalizing on recent advances in quantum field theories, have been exploited to derive an exact formula for the Heun connection matrix and alternative formulations of the waveforms were proposed in terms of the instanton partition function of a $SU(2)^{2}$ quiver gauge theory. Following this path the infinite tower of logarithmic terms (tails and tails of tails) have been resummed into exponentials.
@@
More recently, multipole [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48] and perturbation theory [49] methods have been applied to the study of scattering processes (open trajectories) in gravity [50, 51], a field where we can largely benefit from the impressive developments and refinements of scattering amplitude techniques in the last few years (see [52] for a review). These systems provide a richer set up, where the couplings of the PM and PN expansions are no longer linked to each other. For example, in the scattering amplitude language, the PM expansion corresponds to the expansion in loops. At a given order in the coupling, the amplitude is exact in the velocity, i.e. it resums the infinite PN series. To reproduce this result in perturbation theory, one has to resum the double expansion in one of the two variables.
@@
For simplicity we focus on the geodesic motion in a Schwarzschild geometry and parametrize the mass, wavefrequency and radial distance by the two dimensionless combinations $z=2M/r$ and $y=2i\omega r$, assuming $x=4{\rm i}\omega M$ is small. We will show that, at order $x^{k}$, the dependence of the waveform on the variable $y$ or $z$ can be resummed and written in terms of a single hypergeometric function and its derivative, with coefficients given by two polynomials of degree $k$, that can be determined recursively. The two limits $x\to 0$ and $y$ or $z$ finite lead to two alternative hypergeometric representations of the confluent Heun function, describing the PM and PN approximations respectively. The parameters of the hypergeometric function and the coefficients of the polynomials are all expressed in terms of a characteristic function $a(\ell)$ that, in the MST [10, 11, 12, 13, 14, 15, 16, 17, 18, 19] formulation of the confluent Heun solution, is the ”renormalized angular momentum”, while in the description of this paper [26, 27, 28, 29, 30, 31, 32, 33, 34, 49], it is the Seiberg-Witten (SW) period of the ${\cal N}=2$ supersymmetric gauge theory. In this framework the gauge couplings of the quiver gauge theory are connected to the PM and PN expansion parameters as $q_{1}=z$ and $q_{2}=y$. The same techniques apply to D-branes, topological stars and other gravitational backgrounds governed by an equation of the Heun type. See [53, 54] for a recent application to the case of topological stars.
@@
To keep the material self-contained, in this paper we present a derivation of the Heun solutions from first principles with no reference to supersymmetric gauge theory or localization, which sit in the background as inspiration to some of our ansätze, that nonetheless could be also accepted as such. These results provide a first principle proof of the Heun connection formula derived in [26, 27, 28, 29] from the crossing symmetry of the conformal blocks of a conformal field theory (CFT) . Our results provide an algorithm which is very efficient when implemented in a Mathematica code and that will allow us to compute the gravitational waveform in the case of circular orbits up to the 30th PN order. This latter result will be tested against a numerical solution and the 22nd PN order computation available in literature [20]. Numerical studies of unbounded systems are postponed to future investigations.
@@
This is the plan of the paper: in Section II we derive the PM and PN expansion of the confluent Heun function. In Section III we apply the results of Section II to the study of the gravitational waveforms emitted by a particle moving in a Schwarzschild background. In Section IV we test our results against those obtained via the MST method and numerical integration. In Section V we extend the results to include the case of the non-confluent Heun function. Section VI contains our conclusions.
II The confluent Heun equation
@@
$\beta$ $=$ $-x(\kappa_{1}+\kappa_{3}),\quad\delta=\kappa_{2}+\kappa_{3},\quad\gamma=1-\kappa_{2}+\kappa_{3}$
$\zeta$ $=$ $-x\qquad,\qquad\quad q=u-\kappa_{3}^{2}-x(\kappa_{1}+\kappa_{3})$ (3)
@@
$G_{\alpha}(z)\underset{x,z\to 0\atop x/z\leavevmode\nobreak\ {\rm finite}}{\approx}H^{0}_{\alpha}({\textstyle\frac{x}{z}})\qquad,\qquad G_{\alpha}(z)\underset{x\to 0\atop z\leavevmode\nobreak\ {\rm finite}}{\approx}H^{1}_{\alpha}(z)$ (7)
@@
$H^{0}_{\alpha}(y)$ $=$ $y^{-\alpha a+\kappa_{1}}{}_{1}F_{1}(\kappa_{1}{-}\alpha a,1{-}2\alpha a;y)$
@@
${\rm Near\leavevmode\nobreak\ zone}:G_{\alpha}(z)=z^{\alpha\,a{-}\kappa_{1}}\sum_{i,j=0}^{\infty}G_{\alpha ij}(a)\,z^{i}\,({\textstyle\frac{x}{z}})^{j}$ (11)
with $x\ll z\ll 1$. $a(u)$ given in (9). The quantity $a$ plays the role of a ‘Floquet exponent’, specifying the monodromy of $G_{\alpha}$ around a contour enclosing the points 0 and $x$. As we will see, the solutions of (4) and the connection formulae take a simpler form if written in terms of the variable $a$ rather than $u$. So we write
@@
II.1 PM approximation: exterior region
@@
$P_{0}(y)=1+\sum_{i=1}^{k}\sum_{j=0}^{i-1}c_{ij}\,x^{i}y^{j-i},\widehat{P}_{0}(y)=\sum_{i=1}^{k}\sum_{j=0}^{i-1}\widehat{c}_{ij}\,x^{i}y^{j-i}\$ (15)
are polynomials of order $k$ in $x$ [^1] For $k=1$ M.Billò, M.L.Frau and A.Lerda first obtained (14) resumming the contributions of the Young tableaux entering the partition function of the supersymmetric gauge theory.. We plug this ansätze into (4) and use the differential equation to express higher derivatives of $H^{0}_{\alpha}$ in terms of $H^{0}_{\alpha}$ and its first derivative themselves. Therefore one finds a system of algebraic linear equations that determines the coefficients $c_{ij}$, $\widehat{c}_{ij}$ and $u_{i}$ (whose complete forms are given in (73)). Setting $c_{00}=1$, one finds, for the first term in the recursion
@@
$c_{i,i-1}=u_{i}\,,$
$\gamma_{\kappa}=1+\sum_{i=1}^{\infty}\widehat{c}_{i,i-1}x^{i}={\rm exp}\left\{-\sum_{i=1}^{\infty}{\partial_{\kappa_{1}}u_{i}\over i}x^{i}\right\}\quad$ (17)
@@
The subscript ”inst” reminds the reader that this quantity is interpreted as the instanton prepotential in the gauge theory description of the Heun equation. Indeed in the limit $z\to 0$, where one of the two gauge couplings is turned off, the quiver theory reduces to a ${\cal N}=2$ supersymmetric $SU(2)$ gauge theory with three massive fundamental flavours with prepotential ${\cal F}_{\rm inst}$.
@@
II.2 PN approximation: interior region
@@
Higher orders follow from the ansätze
@@
$P_{1}(z)$ $=$ $1+\sum_{j=1}^{k}\sum_{i=0}^{j-1}d_{ij}\,z^{i-j}x^{j}$
$\widehat{P}_{1}(z)$ $=$ $\sum_{j=1}^{k}\sum_{i=0}^{j}\widehat{d}_{ij}\,z^{i-j}x^{j}$ (26)
@@
$d_{01}=\frac{2\left(a^{2}-\kappa_{1}^{2}\right)}{4a^{2}-1},\quad\widehat{d}_{11}=-\widehat{d}_{01}=\frac{\left(2\kappa_{1}-1\right)}{4a^{2}-1}$ (27)
@@
$\widetilde{H}^{1}_{\alpha}(z)=z^{a-\kappa_{1}}(1{-}z)^{{1{+}\alpha\over 2}(1{-}\kappa_{2}{-}\kappa_{3})}\,_{2}F_{1}\left({\textstyle\frac{1{+}\alpha}{2}}{-}\alpha\kappa_{2}{+}a,{\textstyle\frac{1{+}\alpha}{2}}{-}\alpha\kappa_{3}{+}a;1{+}\alpha\left(1{-}\kappa_{2}{-}\kappa_{3}\right);1{-}z\right)$ (29)
@@
II.3 Boundary conditions and connection formulae
@@
III Black hole perturbation
@@
III.1 Heun gravity correspondence
@@
with $\Delta=r(r-2M)$. We are interested in the $\psi_{4}$ fundamental mode [55, 8] corresponding to $s=-2$. (40) can be put into the CHE form (4) after the identifications (look at Appendix A for details)
@@
$R_{\rm in}(z)$$\approx$ $\left\{\begin{array}[]{lll}B^{\rm in}_{-}e^{-{x\over 2z}}z^{\kappa_{1}-{3\over 2}}+B^{\rm in}_{+}e^{{x\over 2z}}z^{-\kappa_{1}-{1\over 2}}&\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode&\quad z\to 0\\ D^{\rm in}_{-}(1-z)^{1+\kappa_{2}+\kappa_{3}\over 2}&&\quad z\to 1\\ \end{array}\right.$ (44)
$R_{\rm up}(z)$$\approx$ $\left\{\begin{array}[]{lll}B^{\rm up}_{+}e^{{x\over 2z}}z^{-\kappa_{1}-{1\over 2}}&\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode&\quad z\to 0\\ D^{\rm up}_{-}(1-z)^{1+\kappa_{2}+\kappa_{3}\over 2}+D^{\rm up}_{+}(1-z)^{3-\kappa_{2}-\kappa_{3}\over 2}&&\quad z\to 1\\ \end{array}\right.$ (47)
@@
$G(r,r^{\prime})={1\over W}\left\{\begin{array}[]{ccc}R_{\rm in}(r^{\prime})R_{\rm up}(r)&\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\hfil&r^{\prime}<r\\ R_{\rm in}(r)R_{\rm up}(r^{\prime})&\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\hfil&r<r^{\prime}\\ \end{array}\right.$ (48)
@@
$R_{3}(a,y)=\frac{(7-2a)y^{3}}{8(2a-1)(2a-3)(a-1)}+\frac{(11-6a)x}{4(2a-1)}$
$R_{4}(a,y)=\frac{\left(4a^{2}{-}72a{+}163\right)y^{4}}{512(a{-}2)(a{-}1)(2a{-}3)(2a{-}1)}$
$+\frac{\left({-}4a^{2}{+}48a{-}119\right)yx}{64(2a{-}1)(a{-}1)}{+}\frac{\left(8a^{3}{-}4a^{2}{-}18a{+}9\right)x^{2}}{64y^{2}(a{+}1)}$ (56)